OpenStudy (anonymous):

The circumference of a circle is 24 kilometers. What is its radius?Estimate, using 3 for π.

OpenStudy (anonymous):

please give an accurate answer. Thanks

hero (hero):

C = 2πr 24 = 2πr 24/2π = r 12/π = r

hero (hero):

12/3 = r 4 = r